### v1

Abstract
Let $\Gamma$ be a finite set of Jordan curves in the plane. For any curve $\gamma\in\Gamma$, we denote the bounded region enclosed by $\gamma$ as $\tilde{\gamma}$. We say that $\Gamma$ is a non-piercing family if for any two curves $\alpha,\beta\in\Gamma$, $\tilde{\alpha}\setminus\tilde{\beta}$ is a connected region. A non-piercing family of curves generalizes a family of $2$-intersecting curves in which each pair of curves intersect in at most two points. Snoeyink and Hershberger (“Sweeping Arrangements of Curves”, SoCG ’89) proved that if we are given a family $\mathcal{C}$ of $2$-intersecting curves and a fixed curve $C\in\mathcal{C}$, then the arrangement can be swept by $C$, i.e., $C$ can be continuously shrunk to any point $p\in\tilde{C}$ in such a way that the we have a family of $2$-intersecting curves throughout the process. In this paper, we generalize the result of Snoeyink and Hershberger to the setting of non-piercing curves. We show that given an arrangement of non-piercing curves $\Gamma$, and a fixed curve $\gamma\in\Gamma$, the arrangement can be swept by $\gamma$ so that the arrangement remains non-piercing throughout the process. We also give a shorter and simpler proof of the result of Snoeyink and Hershberger and cite applications of their result, where our result leads to a generalization.
keywords:
Sweeping, Pseudodisks, Discrete Geometry, Topology
category:
[…]
1 Introduction
A fundamental algorithmic design technique in Computational Geometry introduced by Bentley and Ottmann [4] is line sweep. In the original context the technique was invented, namely reporting all intersections among $n$ line segments in the plane. The basic idea is to move a vertical line from $x=-\infty$ to $x=+\infty$, stopping at event points (like end-points of segments or intersections of two segments) where updates are made. Since then, the line sweep technique has found a wide variety of applications in Computational Geometry including polygon triangulation, computing Voronoi diagrams and its dual Delaunay triangulations, and computing the volume of a union of regions to name a few. The method can also be generalized to higher dimensions. See the classic books on Computational Geometry [10, 18, 11] for more applications. Besides applications in Computational Geometry, it is also a useful tool in combinatorial proofs. For a simple example, a one dimensional sweep can be used to show that the clique cover number and maximum independent set size in interval graphs are equal. Several other examples appear in Section 3 of this paper. In most applications, the sweep line moves continuously and covers the plane so that each point in the plane is on the sweep line exactly once. There are also applications where instead of sweeping with a line, we sweep with a closed curve like a circle. For instance we can start with an infinite radius circle and shrink it continuously to its center.
Edelsbrunner and Guibas [12] introduced the topological sweep technique and showed the advantage of sweeping with a topological line, i.e., curve that at any point intersects each segment at most once rather than a rigid straight line. They obtained better algorithms to report all intersection points between lines, which by duality improved results known at that time for many problems on point configurations. Chazelle and Edelsbrunner [8] were also able to adapt the topological sweep “in 20 easy pieces” to report the $k$ intersections of $n$ given lines segments in $O(n\log n+k)$ time.
Thus, we can generalize the sweeping technique to sweep a set of pseudolines [^1] A set $\mathcal{L}$ of bi-infinite curves in the plane is a collection of pseudolines the curves in $\mathcal{L}$ pairwise intersect at most once. with a pseudoline. We can generalize further, and do topological sweep with a closed curve. Generalizing from a set of circles in the plane, where any pair of circles intersect in at most two points, a collection of simple closed curves is called an arrangement of pseudocircles iff any pair of curves intersect in at most two points. Note that for any two disks defined by circles or pseudocircles, $D_{1}$, $D_{2}$, the difference $D_{1}\setminus D_{2}$, is a path-connected set.
Showing that it is possible to sweep a collection of pseudolines with one of the pseudolines, and a set of pseudocircles with one of the pseudocircles is much more challenging than sweeping with a line or circle. Snoeyink and Hershberger [23] showed in a celebrated paper that we can sweep an arrangement of 2-intersecting curves with any curve in the family. The result of Snoeyink and Hershberger has found several applications both in Computational Geometry as well as in combinatorial proofs. See for example, [13, 2, 5, 1, 9, 3, 22] and references therein.
It is natural to attempt to generalize the result of Snoeyink and Hershberger to $k$-intersecting curves for $k>2$, i.e., sweep a set of $k$-intersecting curves with a sweep curve that maintains the invariant that the arrangement is $k$-intersecting throughout the sweep. Unfortunately, Snoeyink and Hershberger [23] show that this is not possible for $k>3$, and ask at the end of their paper what intersection property the sweep satisfies if we sweep an arrangement of $k$-intersecting curves. While we do not fully answer their question, we extend their result to show that we can maintain a more general property than 2-intersection while sweeping. Specifically, we show that if the input curves satisfy a topological condition of being non-piercing, then we can sweep with a curve so that the arrangement satisfies this property throughout the sweep. For a Jordan curve $\gamma$, let $\tilde{\gamma}$ denote the bounded region defined by $\gamma$. Two Jordan curves $\alpha,\beta$ are said to be non-piercing if $\tilde{\alpha}\setminus\tilde{\beta}$ is a path connected region, and a family of curves is non-piercing if they are non-piercing pairwise. The main theorem we prove in this paper is the following.
[…]
Let $\Gamma$ be a finite non-piercing family of curves. Given any $\gamma\in\Gamma$ and a point $P\in\tilde{\gamma}$, we can sweep $\Gamma$ with $\gamma$ so that at any point of time during the sweep, the curves remain non-piercing and $P$ remains within $\tilde{\gamma}$
We also give an alternative, and somewhat simpler proof of the fundamental result of Snoeyink and Hershberger. While our result is not significantly simpler, we believe it leads to a cleaner analysis using two conceptual tools, namely minimal lens bypassing and minimal triangle bypassing that may help in other settings involving pseudodisks or non-piercing regions. Further, the proof of Theorem 1 is quite simple compared to the result of Snoeyink and Hershberger, and even our simpler proof of their result (Theorem 7.3). Hence, the main difficulty in sweeping seems to be maintaining the 2-intersection property.
For many algorithmic applications, especially for several packing and covering problems, the restriction that an arrangement is non-piercing is not harder than the corresponding problems for pseudodisks. For example, hypergraphs defined by points and non-piercing regions enjoy a linear shallow-cell complexity [19] and therefore, admit $\epsilon$-nets of linear size and $O(1)$-approximation algorithms for covering problems via quasi-uniform sampling [24, 7]. In particular, it is plausible that the non-piercing condition is the most elementary condition required for which results that work for pseudodisks extend. For example Sariel Har-Peled [15], and Chan and Grant [6] showed that for several covering problems involving simple geometric regions that are not non-piercing, several problems that admit a PTAS for non-piercing regions [19] become APX-hard.
The paper is organized as follows. The notation used in the paper is described in Section 2. We describe applications of Theorem 1 in Section 3. Section 4 describes the sweeping operations and Section 5 describes the basic operations required in the proof. Section 6 contains the proof of Theorem 1, and Section 7 contains a simpler proof of the theorem of Theorem of Snoeyink and Hershberger on 2-intersecting curves.
2 Preliminaries
Let $\vec{S}^{1}$ denote a circle centered at the origin and oriented counter-clockwise around the origin. An oriented Jordan curve $\gamma$ is a continuous injective map from $\vec{S}^{1}\to\mathbb{R}^{2}$ that respects the orientation of $\vec{S}^{1}$.i.e., the origin is mapped to a point in the region bounded by $\gamma$. Unless otherwise stated, by a curve, we will mean an oriented Jordan curve.
In the following, we represent by $\Gamma$ a finite set of oriented Jordan curves. For any $\gamma\in\Gamma$, we denote by $\tilde{\gamma}$, the bounded region defined by $\gamma$. Observe that $\tilde{\gamma}$ is simply-connected [^2] A region $r\subseteq\mathbb{R}^{2}$ is said to be simply-connected if any closed loop can be continuously deformed to a point inside the region..
A traversal of a curve is a walk starting at an arbitrary point on the curve, walking in the direction of its orientation and returning to the starting point. Let $\mathcal{A}(\Gamma)$ denote the arrangement of the curves in $\Gamma$. We assume throughout the paper that the curves are in general position, i.e., no three curves intersect at a point, any pair of curves intersect in a finite number of points, and they intersect transversally (i.e., they cross) at these points. The arrangement consists of vertices (0-dimensional faces), where two curves intersect, arcs (1-dimensional faces) which are segments of curves between vertices and cells (2-dimensional faces) which are the connected regions of $\mathbb{R}^{2}\setminus\Gamma$. The boundary of a cell thus consists of vertices and arcs. Each arc is a portion of a curve, and each vertex is the intersection of two of the curves. The set of curves which contribute one or more arcs to the boundary of a cell are said to define the cell. If a curve $\alpha$ contributes an arc to the boundary of a cell $C$, we say that $C$ lies on $\alpha$.
A digon is a cell $D$ in the arrangement of any two curves $\alpha,\beta\in\Gamma$ so that both $\alpha$ and $\beta$ contribute one arc to the boundary of $D$. Note that $D$ is not necessarily a cell in the arrangement of all curves in $\Gamma$. If a cell in $\mathcal{A}(\Gamma)$ is a digon, we call it a digon cell. A digon defined by curves $\alpha$ and $\beta$ is called a lens if it contained in both $\tilde{\alpha}$ and $\tilde{\beta}$ and it is called a negative lens if it is not contained in either $\tilde{\alpha}$ or $\tilde{\beta}$. A triangle is a cell $T$ in the arrangement of three curves in $\Gamma$ so that each curve contributes one arc to the boundary of $T$. A triangle need not be a cell in $\mathcal{A}(\Gamma)$. If a cell in $\mathcal{A}(\Gamma)$ is a triangle, we call it a triangle cell.
Given an arrangement of oriented Jordan curves $\Gamma$ and a sweeping curve $\gamma\in\Gamma$, and a property $\Pi$ satisfied by the curves in $\Gamma$, and a point $P\in\tilde{\gamma}$, a sweep is a continuous movement of the boundary of $\gamma$ until it shrinks to the point $P$ so that $(i)$ for any two points $t>t^{\prime}$, the curve $\gamma$ at time $t$ is contained in the bounded region of $\gamma$ at time $t^{\prime}$, $(ii)$ the union of the sweep curve at all points in time covers $\tilde{\gamma}$, where $\tilde{\gamma}$ refers to the bounded region of the sweep curve before the start of the sweep, and $(iii)$ at each point in time the arrangement of curves satisfies the property $\Pi$. In particular, if the property $\Pi$ is that the curves are non-piercing, we want to ensure that the curves remain non-piercing throughout the process.
Pseudocircles and Non-piercing regions. If the curves in $\Gamma$ intersect pairwise at at most two points, we call $\Gamma$ a family of pseudocircles and in this case the set $\{\tilde{\gamma}:\gamma\in\Gamma\}$ is called a family of pseudodisks. Let $\alpha$ and $\beta$ be two curves. For $k\in\mathbb{N}$, we say that $\alpha$ and $\beta$ are $k$-intersecting if they intersect in at most $k$ points. We say that $\alpha$ and $\beta$ are non-piercing if $\tilde{\alpha}\setminus\tilde{\beta}$, and $\tilde{\beta}\setminus\tilde{\alpha}$ are path-connected. Otherwise, they are said to be piercing. A finite collection of curves $\Gamma$ is said to be non-piercing if the curves in $\Gamma$ are pairwise non-piercing. If in addition the curves in $\Gamma$ are pairwise $k$-intersecting we say that $\Gamma$ is a $k$-admissible family. Figure 1 shows examples of pseudocircles, a pair of piercing curves, and a pair of non-piercing curves.
(a) Pseudocircles
(b) Non-piercing curves
(c) Piercing curves.
Figure 1: Pseudocircles, non-piercing curves and piercing curves
Let $\alpha$ and $\beta$ be two curves which intersect at $k$ points $x_{1},\cdots,x_{k}$. Let $\sigma_{\alpha}(\beta)$ be the sequence of these intersection points between $\alpha$ and $\beta$ encountered in a traversal of $\alpha$. We define $\sigma_{\beta}(\alpha)$ analogously. We say that the intersections of $\alpha$ and $\beta$ are reverse-cyclic if $\sigma_{\alpha}(\beta)$ and $\sigma_{\beta}(\alpha)$ are the reverse of each other upto cyclic shift.
Lemma 2.1 ([21]).
Two curves $\alpha$ and $\beta$ are non-piercing iff the sequences $\sigma_{\alpha}(\beta)$ and $\sigma_{\beta}(\alpha)$ are reverse-cyclic.
3 Applications
In this section, we describe several applications of Theorem 1. We omit the proofs of some of the theorems in this section since they only require replacing the result of Snoeyink and Hershberger (Theorem 7.3) with our result (Theorem 1).
Every arrangement of non-piercing regions contains a “small” non-piercing region: Pinchasi [17] proved that an arrangement of pseudodisks $\Gamma$ contains a pseudodisk $\gamma$ s.t. $\gamma$ is intersected by at most 156 disjoint pseudodisks in $\Gamma\setminus\{\gamma\}$. Using Theorem 1, Pinchasi’s result extends immediately to non-piercing regions.
Theorem 3.1.
Let $\Gamma$ be an arrangement of non-piercing curves in the plane. Then, there is a curve $\gamma\in\Gamma$ s.t. the number of curves in $\Gamma$ whose regions are disjoint and intersect $\tilde{\gamma}$ is at most 156.
A classic algorithm to compute an independent set in an arrangement $\mathcal{D}$ of disks in the plane is to greedily select a smallest disk and remove the disks intersecting it. For any collection of disks, it is easy to show that there is a disk s.t. there are at most 5 disjoint disks intersecting it. In particular, the disk with smallest radius in $\mathcal{D}$ has this property. This implies that the greedy algorithm described yields a $5$-approximation for the Maximum Independent Set problem in the intersection graph of disks in the plane. From Theorem 3.1, it follows by the same argument that a greedy algorithm yields a 156-approximation for the Maximum Independent Set problem in the intersection graph of non-piercing regions in the plane.
Multi-hitting set with non-piercing regions: Raman and Ray [20] studied Hitting set and Set cover problems with demands for set systems defined by a set of points and set of non-piercing regions in the plane. Their proofs rely on showing the existence of suitable graphs. For one of the results they used the result of Snoeyink and Hershberger and consequently it applied only to pseudodsisks. Using Theorem 1, it can be extended to non-piercing regions.
Theorem 3.2.
Given an instance $(P,\mathcal{S})$ where $P$ is a set of points in the plane and $\mathcal{S}$ is a collection of non-piercing regions in the plane with demands $d:\mathcal{S}\to\mathbb{N}$, there is a PTAS for the Multi-hitting set problem, i.e., selecting the smallest size subset $Q\subseteq P$ s.t. $\forall S\in\mathcal{S}$, $Q\cap S\geq d(S)$ when the demands are bounded above by a constant, and otherwise there is a $(2+\epsilon)$-approximation algorithm.
Number of hyperedges defined by lines and non-piercing regions: Keller et al., [16] considered the hypergraph $(\mathcal{L},\mathcal{D})$, where the elements of the hypergraph are defined by a set $\mathcal{L}$ of lines in the plane, and each pseudodisk $D\in\mathcal{D}$ defines a hyperedge consisting of the lines in $\mathcal{L}$ intersecting a pseudodisk. The authors prove that the number of hyperedges in such a hypergraph is $O(|\mathcal{L}|^{3})$. By using Theorem 1 in place of the result of Snoeyink and Hershberger (Theorem 7.3), and following their result, we obtain the same bound for the hypergraph defined by lines and non-piercing regions.
Theorem 3.3.
Let $(\mathcal{L},\mathcal{D})$ be a hypergraph defined by a set $\mathcal{L}$, where each hyperedge is defined by the set of lines intersecting a non-piercing region. Then, the number of hyperedges of size $t$ in this hypergraph is $O(|\mathcal{L}|^{2})$, and the total number of hyperedges is $O(|\mathcal{L}|^{3})$. Both bounds are tight already for pseudodisks.
Construction of Supports: Raman and Ray[19] proved that any arrangement of non-piercing regions admits a planar support. Using Theorem 1, we obtain a significantly simpler proof.
Theorem 0 (Support for non-piercing regions).
Let $\mathcal{H}$ be a family of non-piercing regions and let $P$ be any set of points in the plane. The hypergraph $(P,\mathcal{H})$ admits a planar support, i.e., there exists a planar graph $G=(P,E)$ s.t. for any $H\in\mathcal{H}$ the subgraph of $G$ induced by $H\cap P$ is connected.
Proof 3.4.
We start with an empty graph $G$ on $P$. For any region $H\in\mathcal{H}$ that contains at least two points from $P$, we add one or more region to the family $\mathcal{H}$ while keeping the family non-piercing as follows. At any point in time let $G_{H}$ denote the graph induced by $G$ on the points in $H$. We repeatedly find a shrunk copy $H^{\prime}$ of $H$ which contains exactly two points of $p$ belonging to different components of $G_{H}$. We add $H^{\prime}$ to $\mathcal{H}$ and add an edge between the two points it contains. We do this until $G_{H}$ is connected. Our next claim shows that such a shrunk copy $H^{\prime}$ exists so that the $\mathcal{H}\cup\{H^{\prime}\}$ remains non-piercing.
Claim 1.
We can find a region $H^{\prime}$ such that i) $\mathcal{H}\cup\{H^{\prime}\}$ is a non-piercing family and ii) $H$ intersects exactly two regions in $\mathcal{K}$ that belong to distinct components of $G_{H}$.
Proof 3.5.
We create a copy of $H$ and call it $H^{\prime}$. We will shrink $H^{\prime}$ to satisfy the constrains in the claim. We first shrink $H^{\prime}$ continuously (using Theorem 1) while it still contains points from at least two of the components of $G_{H}$ stopping right before any further shrinking would cause $H^{\prime}$ to contain points from only one of the components. At this point $H^{\prime}$ contains only one point $p$ from one of the components. It also contains one or more points from another component. We now fix the point $p$ and continue shrinking $H^{\prime}$ using Theorem 1 so that $H^{\prime}$ continues to contain $p$ and we stop when $H^{\prime}$ contains only one point from the other component. At this point, we have the desired $H^{\prime}$.
This finishes our construction of the graph $G$ and by construction, it is a support. We now show that it is planar. Each edge $e$ added to the graph corresponds to a region $H_{e}$ that contains the two end points of $e$. We can now draw the edge $e$ in the plane using any curve joining its two end-points using a curve that lies in $H_{e}$. Now, consider two edges $e$ and $e^{\prime}$. Since their corresponding regions $H_{e}$ and $H_{e^{\prime}}$ are non-piercing, it follows that the drawings of $e$ and $e^{\prime}$ intersect an even number of times. This shows (by the Hanani-Tutte theorem [14]) that the graph $G$ is planar.
4 Sweeping
Let $\Gamma$ be a set of curves in the plane and let $\gamma\in\Gamma$ be one of the curves called the sweep curve. By “sweeping of $\mathcal{A}(\Gamma)$ by $\gamma$”, we mean the process of continuously shrinking $\tilde{\gamma}$ over a finite duration of time until there are no intersection points among the curves in $\Gamma$ in the interior of $\tilde{\gamma}$. In general, we want to do this while maintaining some topological invariants during the sweep. Snoeyink and Hershberger [23] discretized this countinous sweeping process into a set of discrete operations each of which can be implemented as a continuous deformation of $\gamma$ over a unit time interval. We follow their arguments and likewise describe a set of allowable discrete sweeping operations. Most of our operations are slight variations on their operations, except one new operation that we introduce. In order to describe the operations, we need one more definition. We say that a curve $\alpha$ is visible from the sweep curve $\gamma$ if there is an arc $\tau(u,v)$ (called the visibility arc) joining a point $u$ on $\gamma$ with a point $v$ on $\alpha$ whose interior lies in $\tilde{\gamma}$ and does not intersect any of the curves in $\Gamma$. We define the following discrete sweeping operations:
  1. 1.
    Take a loop: Let $\alpha$ be a curve that does not intersect $\gamma$ and such that $\alpha$ is visible from $\gamma$ via the visibility arc $\tau(u,v)$. We define “taking a loop” as the operation that modifies $\gamma$ to $\gamma^{\prime}$ as follows: Taking two points $u^{\prime},u^{\prime\prime}$ very close to $u$ on either side to $u$ on $\gamma$, we replace the segment of $\gamma$ between $u^{\prime}$ and $u^{\prime\prime}$ by a curve between $u^{\prime}$ and $u^{\prime\prime}$ that lies very close to $\tau(u,v)$ and loops around $v$ crossing $\alpha$ twice. See Figures 1(a) and 1(b). More formally, $\gamma$ is modified to $\gamma^{\prime}$ so that $\tilde{\gamma^{\prime}}=\tilde{\gamma}\setminus(\tau(u,v)\oplus B_{\epsilon})$ where $\oplus$ denotes Minkowski sum and $B_{\epsilon}$ is a ball of radius $\epsilon$ for an arbitrarily small $\epsilon$. In the rest of the paper, we avoid such formal definitions and resort to figures for ease of exposition. However, our informal definitions using figures can easily be formalized.
    (a) Visibility curve $\tau(u,v)$ from $\gamma$ to $\alpha$.
    (b) The resulting curve $\gamma^{\prime}$ after taking a loop on $\alpha$.
    Figure 2: Taking a loop.
  2. 2.
    Bypass a digon cell: Suppose that a curve $\alpha$ and the sweep curve $\gamma$ form digon cell $D$ with vertices $u$ and $v$. Then, we define the operation of “bypassing $D$” as the modification of $\gamma$ to a curve $\gamma^{\prime}$ that goes around the digon cell $D$ as shown in Figures 2(a) and 2(b).
    (a) Digon cell $D$ formed by $\alpha$ and $\gamma$.
    (b) The curve $\gamma^{\prime}$ after bypassing $D$.
    Figure 3: Bypassing a digon cell
  3. 3.
    Bypass a triangle cell: Let $\alpha$ and $\beta$ be two curves that along with the sweep curve $\gamma$ define a triangle cell $T$. We define the operation of “bypassing $T$” as the modification of $\gamma$ to a curve $\gamma^{\prime}$ which goes around $T$ as show in Figures 3(a) and 3(b).
    (a) A triangle cell $T$ formed by $\alpha,\beta$ and $\gamma$.
    (b) The resulting curve $\gamma^{\prime}$ after bypassing $T$.
    Figure 4: Bypassing a triangle cell.
  4. 4.
    Bypass a visible vertex: Let $v$ be a vertex lying inside $\tilde{\gamma}$ where the curves $\alpha$ and $\beta$ intersect. If there is a visibility arc $\tau(u,v)$ from a point $u$ on $\gamma$ to $v$ whose interior lies in the interior of $\tilde{\alpha}$ as well as $\tilde{\beta}$, then by definition, the interior of $\tau(u,v)$ also lies in the interior of $\tilde{\gamma}$. In this case, we refer to $\tau(u,v)$ as the bypassability curve for $v$. “Bypassing $v$” is the modification of $\gamma$ to a curve $\gamma^{\prime}$ that takes a loop around $v$ as shown in Figures 4(a) and 4(b). Note that unlike the previous operations, the orientation of the curves are important here.
    (a) Visibility curve $\tau(u,v)$ to $v$, the intersection of $\alpha$ and $\beta$.
    (b) The curve $\gamma^{\prime}$ after bypassing vertex $v$.
    Figure 5: Bypassing a visible vertex.
The following Lemma summarizes properties of the above operations that are intuitively obvious from the figures.
Lemma 2.
Let $\Gamma$ be an arrangement of non-piercing curves with sweep curve $\gamma\in\Gamma$. The operations $(i)$ Take a loop, $(ii)$ Bypass a digon cell, $(iii)$ bypass a triangle cell, and $(iv)$ bypass a visible vertex, leave the arrangement of curves non-piercing but can change the number of intersections between $\gamma$ and other curves. Operation $(i)$ increases the number of intersection points between $\gamma$ and the curve $\alpha$ we take a loop on by $2$. Operation $(ii)$ decreases the number of intersections between $\gamma$ and the other curve $\alpha$ defining the digon cell. Operation $(iii)$ does not change the number of intersections between $\gamma$ and other curves. Finally, operation $(iv)$ increases by $2$ the number of intersections between $\gamma$ and each of the curves $\alpha,\beta$ defining the vertex being bypassed.
Proof 4.1.
Consider operation $(i)$ and let $\alpha$ is the curve on which we are taking a loop. As shown in Figure 2, this increases the number of intersections between $\alpha$ and $\gamma$ by $2$ by inserting two intersection points along either curve. The two new intersections points appear consecutively along either curve but in opposite order along their orientations. This ensures that $\sigma_{\alpha}(\gamma)$ and $\sigma_{\gamma}(\alpha)$ remain reverse cyclic. In other words, $\alpha$ and $\gamma$ remain non-piercing after the operation. Since no other curve is affected, the entire arrangement remains non-piercing.
Now consider operation $(ii)$ and let $\alpha$ be the curve defining the digon cell along with the sweep curve $\gamma$. As Figure 7 shows, in this case we remove two intersections points (namely the vertices of the digon cell) which appear consecutively along both curves. As a result $\sigma_{\alpha}(\gamma)$ and $\sigma_{\gamma}(\alpha)$ remain reverse cyclic i.e., $\alpha$ and $\gamma$ remain non-piercing. No other curves are affected.
Next, consider operation $(iii)$. In this case let $\alpha$ and $\beta$ be the curves defining the triangle along with the sweep curve $\gamma$. As shown in Figure 4, the intersection points of $\gamma$ with $\alpha$ and $\beta$ effective move switching order along $\gamma$. However this does not affect any of the reverse cyclic sequences. Thus the number of intersection points does not change and the curves remain non-piercing.
Finally, consider operation $(iv)$. As shown in Figure 5, the curve $\gamma$ goes around the vertex $v$ being swept in the process intersecting the curves $\alpha$ and $\beta$ defining $v$ two more times each. If we consider the pair of curves $\alpha$ and $\gamma$ and focus on the intersections between them (i.e., we ignore intersections with other curves), the two new intersections are consecutive along both curves. Therefore, $\sigma_{\alpha}(\gamma)$ and $\sigma_{\gamma}(\alpha)$ remain reverse cyclic i.e., $\alpha$ and $\gamma$ remain non-piercing.
Corollary 3.
Let $\Gamma$ be an arrangement of 2-intersecting curves with sweep curve $\gamma\in\Gamma$. Then, applying any of the three operations: taking a loop on a curve that does not intersect $\gamma$, bypassing a digon cell, or bypassing a triangle cell leaves the arrangement two-intersecting.
Proof 4.2.
The only operation that increases the number of intersections is the operation of taking a loop. Since we apply the operation of taking a loop only on a curve that does not intersect $\gamma$, it follows that the arrangement remains 2-intersecting.
Figure 6 shows that in the case of non-piercing curves, unlike the case with 2-intersecting curves, we sometimes require the operation of bypassing a vertex.
Figure 6: An example of non-piercing curves where the only operation available for the sweep is to bypass either the vertex labeled $a$ or the vertex labeled $b$.
5 Minimal Digons and Triangles
In this section, we describe a basic operation required in our proofs, namely that of bypassing minimal digons or minimal triangles in an arrangement $\mathcal{A}(\Gamma)$ of a set of non-piercing curves $\Gamma$. Recall that in general digons or triangles are not cells in the arrangement $\mathcal{A}(\Gamma)$. A digon $L$ formed by curves $\alpha$ and $\beta$ is minimal if it does not contain another digon. Equivalently, $L$ is minimal if the intersection of any curve $\delta$ with $L$, consists of disjoint arcs each having one end-point on $\alpha$ and one end-point on $\beta$. Let $L$ be a minimal digon formed by curves $\alpha$ and $\beta$ with vertices $u$ and $v$. The operation of bypassing $L$ modifies the arcs of $\alpha$ and $\beta$ as shown in Figure 7. Formally, the bypassing is done in two steps. First we replace the arc of $\alpha$ between $u$ and $v$ by the arc of $\beta$ between $u$ and $v$ and vice versa so that $u$ and $v$ are now points of tangencies between $\alpha$ and $\beta$. Next we get rid of the tangencies by moving $\alpha$ and $\beta$ slightly apart from each other around $u$ and $v$. We also add reference points called $u_{\alpha}$ and $u_{\beta}$ arbitrarily close to $u$ so that $\alpha$ passes between $u$ and $u_{\alpha}$ keeping $u_{\beta}$ on the same side as $u$, and similarly $\beta$ passes between $u$ and $u_{\beta}$ keeping $u_{\alpha}$ on the same side as $u$. We define the reference points $v_{\alpha}$ and $v_{\beta}$ close to $v$ analogously. Note that the reference points are not vertices in the modified arrangment. We claim that the if a minimal digon in a non-piercing family is bypassed, the modified family obtained is still non-piercing.
Figure 7: The operation of digon bypassing of curves $\alpha$ and $\beta$. The dotted segments are the intersection of a single curve $\delta\in\Gamma$ with the lens $L$.
Lemma 4.
Let $L$ be a minimal digon formed by $\alpha$ and $\beta$ in the arrangement $\mathcal{A}(\Gamma)$ as above. The modified family of curves obtained by bypassing $L$ is non-piercing. Furthermore, the number of intersections between $\alpha$ and $\beta$ decreases by $2$ and for any other pair of curves the number of intersections does not change.
Proof 5.1.
Since $L$ is a minimal digon, the intersection of any other curve $\delta\in\Gamma\setminus\{\alpha,\beta\}$ is a collection of disjoint segments with one vertex on $\alpha$ and the other on $\beta$. Let $s$ be one of these segments with end point $x$ on $\alpha$ and $x^{\prime}$ on $\beta$. Then, as a result of bypassing, the earlier intersection of $\alpha$ and $\delta$ at $x$ is replaced by the intersection of the modified $\alpha$ and $\delta$ at $x^{\prime}$. However, since $\alpha$ and $\delta$ do not have any intersections in the interior of the segment $s$, the order of the intersections of $\alpha$ and $\delta$ along either curve remains unchanged i.e., they remain reverse cyclic. Hence, by Lemma 2.1, the modified $\alpha$ and $\delta$ are non-piercing. Analogously the modified $\beta$ and $\delta$ are non-piercing. This also shows that the number of intersection between $\alpha$ and $\delta$ does not change. Similarly the number of intersections between $\beta$ and $\delta$ does not change. The curves $\alpha$ and $\beta$ lose two consecutive points of intersections (namely the vertices of $L$) and therefore the order of the remaining intersections along them remains reverse cyclic. Thus, the modified $\alpha$ and the modified $\beta$ are also non-piercing.
The only new/changed cells in the modified arrangement obtained after bypassing $L$ are those that contain one of the points $u,v,u_{\alpha},u_{\beta},v_{\alpha}$ or $v_{\beta}$. Note that there may also be cells that are identical to an old cell but their defining set of curves has changed - $\alpha$ is replaced by $\beta$ or vice versa.
Lemma 5.
Only the new cells containing one of the reference points $u_{\alpha}$, $u_{\beta}$, $v_{\alpha}$ or $v_{\beta}$ may be newly created digon/triangle cells on the sweep curve $\gamma$ in the arrangement obtained after bypassing $L$.
Proof 5.2.
The number of sides in the cells other than those containing the reference points $u_{\alpha},u_{\beta},v_{\alpha}$ or $v_{\beta}$, or the points $u$ or $v$ are identical to an old cell. One of their sides defined by $\alpha$ may have been replaced by $\beta$ or vice versa but this does not change whether the cell lies on $\gamma$. Thus such cells cannot be digon/triangle cells on $\gamma$ unless they were already so in the original arrangement (before bypassing $L$). The cell in the new arrangement containing $u$ an
Let $T$ be a triangle on the sweep curve $\gamma$ defined by the curves $\alpha,\beta\in\Gamma$ along with $\gamma$. We say that $T$ is a “minimal triangle on $\gamma$”, or “base $\gamma$” if the intersection of any other curve $\delta\in\Gamma\setminus\{\alpha,\beta,\gamma\}$ with $T$ consists of a set of disjoint arcs each with one end-point on $\alpha$ and one end-point on $\beta$. The operation of bypassing $T$ modifies $\alpha$ and $\beta$ so that they go around $T$ and their intersection $u$ on the boundary of $T$ moves outside $\tilde{\gamma}$ as shown in Figure 8. We show next, that bypassing $T$ leaves the arrangement non-piercing.
Figure 8: The operation of minimal triangle bypassing of curves $\alpha$ and $\beta$. The dotted segments are the intersection of a curve $\delta\in\Gamma$ with the triangle $T$.
Lemma 6.
Let $T$ be a minimal triangle in an arrangement $\Gamma$ of non-piercing regions defined by curves $\alpha,\beta$ and $\gamma$. Then, bypassing $T$ yields a non-piercing arrangement. The number of intersection points between any pair of curves does not change. However, the number of intersection points in $\tilde{\gamma}$ decreases by 1. Thus, the number of intersections in $\tilde{\gamma}$ decreases by 1.
Proof 5.3.
Let $u$ be the intersection of $\alpha$ and $\beta$ on the boundary of $T$. Similarly, let $w$ be the intersection of $\alpha$ and $\gamma$ on the boundary of $T$.
By the definition of the operation of bypassing $T$ (shown in Figure 8), the intersection $u$ between $\alpha$ and $\beta$ that lies in $\tilde{\gamma}$ is replaced by a new intersection $u^{\prime}$ that lies outside $\tilde{\gamma}$. The following arguments also show that the number of intersections between any other pair of curves within $\tilde{\gamma}$ does not change. Thus, the total number of intersections within $\tilde{\gamma}$ decreases by $1$ when $T$ is bypassed.
We now argue that all pairs of curves remain non-piercing and the number of pairwise intersections remains unchanged except for the pair $\alpha,\beta$. Let $\delta$ be any curve in $\Gamma\setminus\{\alpha\}$. We show that the modified $\alpha$ and $\delta$ are non-piercing and when $\delta\neq\beta$ the number of intersection $\alpha$ and $\delta$ within $\gamma$ remains the same. An analogous statement holds for the modified $\beta$ and $\delta$ for any $\delta\in\Gamma\setminus\{\beta\}$.
Case 1 : $\delta=\gamma$. The modified $\alpha$ and $\gamma$ intersect at a new point $v$ instead of at $w$ (see Figure 9). Since $\alpha$ and $\gamma$ don’t intersect at any point between $v$ and $w$ on the boundary of $T$, their reverse cyclic sequences remain the same and therefore they remain non-piercing.
Case 2: $\delta=\beta$. The modified $\alpha$ and modified $\beta$ intersect at $u^{\prime}$ instead of $u$. Since $\alpha$ and $\beta$ did not have any intersections other than $u$ on the boundary of $T$, this does not change their reverse cyclic sequences. Thus, they remain non-piercing.
Case 3: $\delta\in\Gamma\setminus\{\alpha,\beta,\gamma\}$. If $\delta$ does not intersect $T$, the intersection points between $\alpha$ and $\delta$ does not change and therefore they remain non-piercing. Let us suppose therefore that $\delta$ intersects $T$. Its intersection with $T$ then consists of a disjoint non-intersecting set of arcs with one end-point each on the two sides of $T$ defined by $\alpha$ and $\beta$. Bypassing $T$ moves the intersection between $\delta$ and $\alpha$ from one end-point to the other on each of the arcs. However since the arcs are non-intersecting, this does not affect their reverse cyclic sequences. They thus remain non-piercing.
Corollary 5.4.
Let $T$ be a minimal triangle in an arrangement $\Gamma$ of 2-intersecting curves defined by $\alpha,\beta$ and $\gamma$. Then, bypassing $T$ yields a 2-intersecting arrangement.
Let $T$ be a minimal triangle on $\gamma$ defined by $\alpha$ and $\beta$. Let $u,v$ and $w$ be the vertices of $T$ defined by the pairs $(\gamma,\alpha)$, $(\alpha,\beta)$, and $(\beta,\gamma)$ respectively. Note that $u$ is a vertex of a unique lens defined by $\beta$ and $\gamma$. Let us call the other vertex of this lens $u^{\prime}$. We define $w^{\prime}$ similarly so that $w$ and $w^{\prime}$ are vertices of a lens defined by $\alpha$ and $\gamma$.
Figure 9: The operation of minimal triangle bypassing of curves $\alpha$ and $\beta$.
Lemma 7.
If there is at least one other curve $\delta$ that intersects the minimal triangle $T$, then only the cells containing the reference points $v_{\alpha}$ or $v_{\beta}$ may be digon/triangle cells.
Proof 5.5.
The only cells in the modified arrangement that were not already present in the original arrangement are the ones that contain the points $v,v_{\alpha}$ and $v_{\beta}$. The cell containing $v$ cannot be a digon/triangle cell on $\gamma$ since it has at least four sides namely $\alpha,\beta,\gamma$ and an additional curve that crosses the triangle $T$.
6 Sweeping Non-piercing regions
See 1
Proof 6.1.
Let us call a sweeping operation valid if it does not move $P$ out of $\tilde{\gamma}$. Note that vertex bypassing operations are always valid since we can implement them in such a way that $P$ in not moved out of $\tilde{\gamma}$. The only possible sweeping operations that may move $P$ out of $\tilde{\gamma}$ are therefore digon/triangle bypassing.
If we can always find a valid sweeping operation, we can continue applying them eventually ending up in a situation where no pair of curves intersects in the interior of $\tilde{\gamma}$. By Lemma 2, each sweeping operation ensures that the curves are non-piercing. Thus, for contradiction assume that there exists an arrangement of curves so that there are no valid sweep operations. Among such examples let $\Gamma$ be the simplest in the following sense: it lexicographically minimizes $(m,n)$ where $m$ is the number of curves in $\Gamma$ that lie in the interior of $\tilde{\gamma}$ and $n$ is the number of vertices in $\mathcal{A}(\Gamma)$ that lie in $\tilde{\gamma}$ (including those that lie on $\gamma$). If no pair of curves in $\Gamma$ intersects more than twice in $\tilde{\gamma}$, then by the result of Snoeyink et al., (Theorem 7.3), there is at least one valid sweeping operation.
The result of Snoeyink et al., can also be applied if some curve $\alpha$ intersects $\gamma$ in more than $2$ points, but no pair of curves intersects more than twice in $\tilde{\gamma}$, since we can treat each segment of $\alpha$ in $\tilde{\gamma}$ as a distinct curve and apply their result. We may thus assume that no curve intersects $\gamma$ in more than $2$ points. Since we assumed that no sweep operations are valid, there must be a pair of curves in $\Gamma\setminus\{\gamma\}$ that intersect $3$ or more times in $\tilde{\gamma}$ and therefore form a negative lens inside $\tilde{\gamma}$. Recall that a negative lens is a digon that is not contained in regions bounded by the curves defining it. Let $L$ be a minimal negative lens (i.e., it does not contain another negative lens) in $\tilde{\gamma}$ visible from $\gamma$, which is defined by the curves $\alpha$ and $\beta$. By minimality of $\Gamma$, such a lens should exist. Let $u$ and $v$ be the two intersection points of $\alpha$ and $\beta$ on the boundary of $L$ so that $\alpha$ is oriented from $u$ to $v$ and $\beta$ is oriented from $v$ to $u$. Suppose now that we bypass $L$. The vertices $u$ and $v$ are lost as a result of the bypassing of $L$. We claim that no vertex becomes bypassable as a result of bypassing $L$. The only vertices that can potentially become bypassable are those that lie in $L$. Let $x_{\alpha}$ be a vertex on the boundary of $L$ and that becomes bypassable. We assume without loss of generality that it lies on the modified curve $\alpha$. Its bypassibility curve then arrives at $x_{\alpha}$ from the left of the modified curve $\alpha$ (i.e., from the interior of the new $\tilde{\alpha}$) which means that such a curve must pass arbitrarily close to either $u$ or $v$. Assume that it passes close to $u$ (see Figure 9(a)), the other case being analogous. We can then modify it to terminate at $u$ (as shown in Figure 9(b)) so that it arrives at $u$ from the left of both $\alpha$ and $\beta$ (i.e., from the interior of both $\tilde{\alpha}$ and $\tilde{\beta}$) in the original arrangement. This shows that $u$ was bypassable before we bypassed the lens $L$ which by assumption was not the case. Similarly the vertices on the curve $\beta$ cannot become bypassable as a result of bypassing $L$. An analogous argument also shows that vertices lying in the interior of $L$ (in the original arrangement) cannot become bypassable after $L$ is bypassed.
(a) Bypassability curve for $x_{\alpha}$ in the modified arrangement
(b) Bypassability curve for $u$ in the original arrangement
Figure 10: If a vertex becomes bypassable after we bypass $L$, the corresponding bypassability curve can be modified to a bypassability curve for $u$ in the original arrangement.
Figure 11: If any of the cells containing $u$ or $u_{\alpha}$ is a digon/triangle cell after bypassing the lens $L$ then $u$ must have been bypassable in the original arrangement.
We now claim that the new cells created as a result of bypassing $L$ cannot be digon/triangle cells on $\gamma$. The only cells that are could potentially be digon/triangle cells after bypassing $L$ are those that contain one of the points $u,v,u_{\alpha},u_{\beta},v_{\alpha}$ or $v_{\beta}$. The other cells may have $\alpha$ replaced by $\beta$ (or vice versa) on their boundary as result of bypassing $L$ but this does not change their shape or whether they are on $\gamma$. Since they were not digon/triangle cells on $\gamma$ before, they are not digon/triangle cells after bypassing $L$.
The cells containing $u$ or $u_{\alpha}$ cannot be a digon/triangle cell on $\gamma$ cell since either of those situations would require that the point $p$ shown in Figure 11 (formally $p$ is the first intersection point of $\alpha$ and $\gamma$ that we arrive at if we start at $u$ on the original curve $\alpha$ and walk along it in the direction opposite to its orientation) is a vertex of the digon/triangle cell. However, that implies that in the original arrangement, we could construct a bypassability curve for $u$ as follows: start at a point on $\gamma$ arbitrarily close to $p$ that lies to the left of $\alpha$ and follow $\alpha$ closely until $u$ and terminate there - thus arriving at $u$ from the left of both $\alpha$ and $\beta$ (see Figure 11). This implies that $u$ was bypassable in the original arrangement contradicting our assumptions. By symmetry, the cells containing any of the other points $u_{\beta},v,v_{\alpha}$ and $v_{\beta}$ also cannot be digon/triangle cells. This shows that $u$ must have been bypassable before we bypassed $L$, contradicting our assumption that this was not the case.
Remark: While our result states that we can sweep by continuously shrinking $\gamma$ to a specified point $P$ in $\tilde{\gamma}$, the proof can be easily modified to show that we can sweep by expanding $\gamma$ so that a specified point $Q$ remains outside $\tilde{\gamma}$. If $Q$ is chosen to be a point at infinity, then this shows that we can sweep by expanding $\gamma$ to infinity retaining the property that the curves are non-piercing. The only change required in the proof is that a vertex $q$ defined by curves $\alpha$ and $\beta$ is said to be visible if there is a visibility curve $\tau$ from a point on $\gamma$ to $q$ s.t. the interior of $\tau$ lies outside $\tilde{\gamma},\tilde{\alpha}$, and $\tilde{\beta}$.
7 A simpler proof of the theorem of Snoeyink and Hershberger
In this section, we give a shorter alternative proof of the result of Snoeyink and Hershberger stated in Theorem 7.3. While our proof follows their general framework, we partition into cases differently and the analysis of the cases is relatively simpler.
Let $\Gamma$ be an arrangement of 2-intersecting curves with sweep curve $\gamma$, and let $P$ be a specified point in $\tilde{\gamma}$. Our goal is to sweep the arrangement of $\Gamma$ with $\gamma$ so that the arrangement remains 2-intersecting throughout the process and $P$ remains inside $\tilde{\gamma}$ until no curve in $\Gamma\setminus\gamma$ intersects $\tilde{\gamma}$. We do not use the operation of bypassing a visible vertex, as this violates the 2-intersecting property of the curves. We restrict ourselves to the remaining sweeping operations, namely: $(i)$ taking a loop on a curve that lies entirely in $\tilde{\gamma}$, $(ii)$ bypassing a digon cell, and $(iii)$ bypassing a triangle cell.
A sweeping operation is said to be valid if it does not move $P$ out of $\tilde{\gamma}$. A cell in the arrangement is said to be sweepable if it is a digon/triangle cell on $\gamma$ that does not contain $P$. Otherwise, it is non-sweepable. If at any point in time, there is a sweepable cell, i.e., there is a valid sweeping operation, we say that the arrangement is sweepable. Otherwise, we say that it is non-sweepable.
We will show that the arrangement of any set of 2-intersecting curves with a sweep curve $\gamma$ and a point $P\in\tilde{\gamma}$ is sweepable. For contradiction, assume that this is not true, and among arrangements that are non-sweepable consider the simplest one in the following sense: it lexicographically minimizes the tuple $(c,n,\ell)$ where $c$ is the number of curves in $\Gamma$ lying entirely in the interior of $\tilde{\gamma}$, $n=|\Gamma|$ and $\ell$ is the number of intersection points (among curves in $\Gamma$) lying in $\tilde{\gamma}$ (including those on $\gamma$).
Broad idea. In the restricted setting where each pair of curves in $\Gamma\setminus\{\gamma\}$ intersects at most once inside $\tilde{\gamma}$, we can show that the arrangement is sweepable. Therefore, a simplest non-sweepable arrangement must have at least one pair of curves that intersect twice (i.e., they form a digon) in $\tilde{\gamma}$. In this case, we carefully modify the arrangement by bypassing a minimal digon or a minimal triangle (defined in Section 5) to obtain a simpler non-sweepable arrangement, thus arriving at a contradiction.
We start with the following definition that is borrowed from the paper of Snoeyink and Hershberger [23]. Let $T$ be a triangle defined by curves $\alpha$ and $\beta$ on $\gamma$. We say that $T$ is a half-triangle with edge $\beta$ if the side of the triangle defined by $\beta$ does not intersect any curve in $\Gamma$ in its interior. Figure 12 shows a half-triangle with edge $\beta$.
Figure 12: The figure shows a half-triangle $T$ with edge $\beta$ containing a triangle cell $T^{\prime}$.
Lemma 8.
Let $\Gamma$ be a set of 2-intersecting curves with sweep curve $\gamma$ s.t. all curves in $\Gamma\setminus\{\gamma\}$ intersect $\gamma$, and any pair of curves in $\Gamma\setminus\{\gamma\}$ intersect at most once in $\tilde{\gamma}$. Then, there is a sweepable cell on $\gamma$, i.e., there is either a digon/triangle cell on $\gamma$ that does not contain $P$.
In order to prove the above lemma, we need the following claim which though not explicitly stated is also proved as part of Lemma 3.1 of [23].
Claim 9.
If curves $\alpha$ and $\beta$ form a half-triangle $T$ on $\gamma$ with edge $\alpha$, so that each curve in $\Gamma$ has at most one intersection with the interior of the side of $T$ on $\gamma$, then there is a triangle cell $T^{\prime}$ on $\gamma$ in $T$.
Proof 7.1.
We prove by induction on the number of curves intersecting the side of $T$ on $\gamma$. If no curve intersects the interior of this side, since the curves are 1-intersecting in $\tilde{\gamma}$, it follows that $T$ is a triangle-cell. Assume that the claim holds for any half-triangle with less than $k$ curves (where $k\geq 1$) intersecting the side of $T$ on $\gamma$. Suppose now that there are $k$ curves intersecting the interior of the side of $T$ on $\gamma$.
Let $x$ denote the intersection point of $\beta$ and $\gamma$ on the boundary of $T$. Let $v$ denote the intersection point of $\alpha$ and $\beta$ on the boundary of $T$. For every curve $\delta^{\prime}$ intersecting the interior of the side $T$ on $\gamma$, $\delta^{\prime}$ intersects it at one point, and $\delta^{\prime}$ and $\gamma$ intersect at two points. Therefore, since $T$ is a half-triangle with edge $\alpha$, $\delta^{\prime}$ intersects the side $T$ on $\beta$. Walking from $x$ to $v$ along $\beta$, let $\delta$ be the first curve intersecting $\beta$. The curve $\delta$ intersects the interior of the side of $T$ on $\gamma$. Now, $\beta$ and $\delta$ form a half-triangle $T^{\prime\prime}$ on $\gamma$ with edge $\beta$ and s.t. less than $k$ curves intersect the interior of the side of $T^{\prime\prime}$ on $\gamma$. Since every curve in $\Gamma$ has at most one intersection point on the side of $T$ on $\gamma$, the same holds for $T^{\prime\prime}$. Hence, by the inductive hypothesis, there is a triangle cell $T^{\prime}$ on $\gamma$ that lies in $T^{\prime\prime}$. Since $T^{\prime\prime}$ lies in $T$, $T^{\prime}$ is the claimed triangle cell in $T$.
Proof 7.2 (Proof of Lemma 8).
Each curve $\alpha\in\Gamma$ has two intersection points on $\gamma$ which split $\gamma$ into two arcs. One of these arcs along with $\alpha$ bounds a portion of $\tilde{\gamma}$ that does not contain $P$. We denote this arc by $I_{\alpha}$. The containment order on the arcs induces a partial order $\prec$ on the curves in $\Gamma\setminus\{\gamma\}$ ($\alpha\prec\beta\Leftrightarrow I_{\alpha}\subseteq I_{\beta}$). Let $\alpha$ be a minimal curve with respect to $\prec$. If the digon $D$ defined by $\alpha$ and $\gamma$ is a digon cell, then we are done. Otherwise, let $i$ and $j$ be the vertices of $D$ so that $I_{\alpha}$ is the arc from $i$ to $j$ along $\gamma$ in the direction of its orientation. Let $\beta$ be the first curve intersecting $\alpha$ when following the arc of $\alpha$ on the boundary of $D$ from $i$ to $j$. Since each curve in $\Gamma\setminus\{\gamma\}$ intersects $\gamma$ twice, and pairwise intersect at most once in $\tilde{\gamma}$, $\beta$ intersects both $\alpha$ and $\gamma$ exactly once on the boundary of $D$. Thus, $\alpha,\beta$ and $\gamma$ form a half-triangle $T$ with edge $\alpha$. Note that $T$ lies within $D$, and therefore does not contain $P$. By Claim 9, $T$ contains a triangle cell, and this completes the proof.
The proof of the main result in this section (Theorem 7.3) follows the structure of Theorem 1, i.e., proof for non-piercing regions. Unlike in the non-piercing case however, since the curves are 2-intersecting, no pair of curves form a negative lens. Let $L$ be a minimal lens in a simplest non-sweepable arrangement $\Gamma$ defined by curves $\alpha,\beta\in\Gamma$. We split the proof into two cases: In the first case, there is a curve $\delta\in\Gamma\setminus\{\alpha,\beta\}$ intersecting $L$. In this case, we show that we can modify the arrangement to obtain a simpler non-sweepable arrangement, thus arriving at a contradiction. If no curve in $\Gamma\setminus\{\alpha,\beta\}$ intersects $L$, then we split the proof into further sub-cases: Since $\Gamma$ was assumed to be a simplest non-sweepable arrangement, bypassing $L$ results in a sweepable arrangement. Thus, there is either a digon/triangle cell on $\gamma$. If the a newly created cell is a digon cell on $\gamma$, we show that removing the curve forming the digon cell results in a simpler non-sweepable arrangement. If we create a triangle cell, we require a slightly more elaborate modification to obtain a simpler non-sweepable arrangement. Thus, in all cases, we obtain a contradiction, and hence there is always either a digon/triangle cell on $\gamma$.
Theorem 7.3 ([23]).
Given any family of 2-intersecting curves $\Gamma$, a curve $\gamma\in\Gamma$ and a point $P\in\tilde{\gamma}$, $\Gamma$ can be swept by $\gamma$ using the following operations: i) passing a digon cell, ii) passing a triangle cell and iii) taking a loop, so that at any point in time during the sweep, the family of curves remain 2-intersecting, and $P$ lies in the interior of $\tilde{\gamma}$.
Proof 7.4.
Suppose the statement is not true. Consider a simplest non-sweepable family $\Gamma$ with sweep curve $\gamma$ i.e., the simplest family for which the statement does not hold. Every curve in this family must intersect $\tilde{\gamma}$ as otherwise we can remove it from $\Gamma$ to obtain a simpler non-sweepable family. If there is a curve $\alpha$ lying entirely in $\tilde{\gamma}$ it cannot be visible from $\gamma$, as otherwise we could take a loop on $\alpha$ (while still keeping $P$ within $\tilde{\gamma}$) to obtain a simpler non-sweepable family - since this decreases the number of regions contained in $\gamma$. If $\alpha$ is not visible from $\gamma$, $\Gamma^{\prime}=\Gamma\setminus\{\alpha\}$ is a simpler non-sweepable arrangement contradicting the minimality of $\Gamma$. If the curves in $\Gamma\setminus\{\gamma\}$ pairwise intersect at most once in $\tilde{\gamma}$, then by Lemma 8, the arrangement is sweepable.
Therefore, in the arrangement $\Gamma$, we can assume that $(i)$ each curve in $\Gamma\setminus\{\gamma\}$ intersects $\gamma$ twice, and there is a pair of curves intersecting twice in $\tilde{\gamma}$, i.e., they form a digon in $\tilde{\gamma}$, $(ii)$ each minimal lens in $\tilde{\gamma}$ is visible from $\gamma$, and $(iii)$ there is either no digon/triangle cell on $\gamma$, or there is exactly one digon/triangle $C$ cell in the arrangement, and $C$ contains the point $P$.
Let $L$ be a minimal digon contained in $\tilde{\gamma}$ and suppose that $L$ is defined by the curves $\alpha,\beta\in\Gamma\setminus\{\gamma\}$ and has vertices $u$ and $v$. $L$ must be visible from $\gamma$, as otherwise by Lemma 4, bypassing $L$ results in a simpler non-sweepable arrangement. Let $a_{1}$ and $a_{2}$ be the intersection points of $\alpha$ with $\gamma$ s.t. the points $a_{1},u,v,a_{2}$ lie in cyclic order along $\alpha$.Similarly, let $b_{1}$ and $b_{2}$ be the intersection points of $\beta$ and $\gamma$ so that $b_{1},u,v,b_{2}$ appear in cyclic order along $\beta$. See Figure 13. Let $u_{\alpha},u_{\beta},v_{\alpha}$ and $v_{\beta}$ be the reference points for bypassing lens $L$. We now split the proof into two cases depending on whether $L$ is non-empty, i.e., there is a curve $\delta\in\Gamma\setminus\{\alpha,\beta,\gamma\}$ that intersects $L$, or is empty, i.e., there is no such curve.
Case 1: $L$ is non-empty.
(a) A minimal digon that is non-empty.
(b) A minimal digon that is empty.
Figure 13: The two cases in the proof. Either $L$ is empty or non-empty.
(a) A triangle $T$ formed on bypassing $L$.
(b) The minimal triangle shaded.
(c) Bypassing the minimal triangle.
Figure 14: Bypassing the minimal triangle.
The proof of this case requires the following technical lemma.
Claim 10.
There exists a curve $\delta\in\Gamma$ that intersects $L$ and such that intersections of any of the other curves with $\delta$ within $\tilde{\gamma}$ lie in $L$. In other words, there are no intersection points on $\delta$ that lie in $\tilde{\gamma}$ but outside $L$.
Proof 7.5.
Bypassing $L$ should create a sweepable cell as otherwise the simpler arrangement obtained on bypassing $L$ remains non-sweepable contradicting the minimality of $\Gamma$. By Lemma 5, the sweepable cell created must contain one of the reference points $u_{\alpha},u_{\beta},v_{\alpha}$ or $v_{\beta}$. By symmetry, we can assume without loss of generality that it contains $u_{\alpha}$.
Consider the initial arrangement where $L$ has not been bypassed as in Figure 12(a). Since there is at least one curve intersecting $L$, we pick $\delta$ as the curve that has the first intersection $y$ with $\beta$ as we walk on $\beta$ from $u$ to $v$. The intersection of $\delta$ with $L$ consists of disjoint segments with end-points on the boundary of $L$. Since $L$ is a minimal lens, each such segment has one end-point on $\alpha$ and one on $\beta$.
Let $x$ be the other end-point of the segment whose one end-point is $y$. Note that $x$ lies on $\alpha$. Let $d_{1},d_{2}$ be the intersections of $\delta$ and $\gamma$ so that $d_{1},x,y,d_{2}$ appear in cyclic order along $\delta$. The cell $C$ containing $u_{\alpha}$ after bypassing $L$ cannot be a digon cell on $\gamma$ since such a digon cell must be the digon defined by $\alpha$ and $\gamma$ and containing $u_{\alpha}$. However that digon contains the point $y$ and cannot be a cell in the arrangement. $C$ can however be a triangle cell on $\gamma$ and in this case it is defined by the curves $\alpha,\gamma$ and $\delta$ and has vertices $a_{1},y$ and $d_{2}$ (Figure 13(a)). Note that this requires that $d_{2}$ lies on the arc of $\gamma$ between $a_{1}$ and $a_{2}$. Also note that if this is the case then, in the original arrangement (before bypassing $L$), there are no intersection points between $a_{1}$ and $u$ on $\alpha$ and between $u$ and $y$ on $\beta$. This along with the fact that $L$ is minimal implies that triangle $T$ with vertices $x,a_{1}$ and $d_{2}$ is a minimal triangle on $\gamma$. Therefore by Lemma 6 we can bypass it (as shown in Figure 13(b)) so that the resulting arrangement is still $2$-intersecting and has one less vertex (since we lose the intersection $x$) inside $\tilde{\gamma}$. Since the new arrangement is simpler, it must be sweepable which means there must be a digon/triangle cell on $\gamma$ not containing $P$ created as a result of bypassing $T$. By Lemma 7, only the cells containing the reference points $x_{\alpha}$ or $x_{\delta}$ can possibly be a digon/triangle cell on $\gamma$. However, the cell containing $x_{\alpha}$ cannot be on $\gamma$ as it is contained in a lens formed by the modified $\alpha$ and $\beta$ that lies in the interior of ${\gamma}$. Therefore, only the cell containing $x_{\delta}$ could be a digon/triangle cell on $\gamma$. However, this cell can lie on $\gamma$ only if $d_{1}$ lies in the arc of $\gamma$ between $b_{2}$ and $b_{1}$ (see Figure 13(c)). In particular, this implies that the interior of the arc of $\delta$ between $x$ and $d_{1}$ is not intersected by any curve in the original arrangement $\Gamma$. Thus, $\delta$ is the desired curve.
Since $\Gamma$ was not sweepable, all cells on $\gamma$ excluding the cell containing $P$ have at least 4 vertices. Now, consider the arrangement $\Gamma^{\prime}=\Gamma\setminus\{\delta\}$, where $\delta$ is the curve guaranteed by Claim 10. We claim that Γsuperscriptnormal-$\Gamma^{\prime}$ remains non-sweepable. Note that the arc of $\delta$ between $d_{2}$ and $y$ separates two cells $C_{1}$ and $C_{2}$. Either one of them contains $P$, or both cells have at least four vertices. Thus, the new cell created by merging these two cells upon removal of $\delta$ is non-sweepable. An analogous statement holds for the arc of $\delta$ between $d_{1}$ and $x$. Finally note that the remaining new cells created as a result of removing the portion of $\delta$ between $x$ and $y$ are not on cells on $\gamma$ since they lie within the lens $L$. Thus, Γsuperscriptnormal-$\Gamma^{\prime}$ is a simpler non-sweepable arrangement contradicting the minimality of $\Gamma$.
Case 2. $L$ is empty. In this case again, we obtain a contradiction by finding a simpler non-sweepable arrangement. Since we assumed that $\Gamma$ is a simplest non-sweepable arrangement, bypassing $L$ must create a sweepable cell on $\gamma$.
Let $D_{u}$ be the cell in $\Gamma$ contained in $\tilde{\gamma}$ with $u$ as a vertex, and has portions of the arcs $[a_{1},u]$ and $[b_{1},u]$ as two sides. Similarly, let $D_{v}$ be the cell containing vertex $v$, and portions of the arcs $[a_{2},v]$ and $[b_{2},v]$ as sides. By Lemma 5, the only digon/triangle cells on $\gamma$ in the new arrangement can be the cells that contain one of the reference points $u_{\alpha},u_{\beta},v_{\alpha}$, or $v_{\beta}$ (See Figure 7). Since $L$ was empty, there is a single cell containing $u_{\alpha}$ and $v_{\alpha}$. Let $C_{\alpha}$ denote this cell. Similarly, let $C_{\beta}$ denote the cell containing the reference points $u_{\beta}$ and $v_{\beta}$.
We treat $C_{\alpha}$ and $C_{\beta}$ independently. Let $C$ denote one of these regions. If $C$ is non-sweepable, we don’t do anything. If $C$ is a digon cell not containing $P$, we throw away the corresponding curve ($\alpha$ or $\beta$). If $C$ is a triangle cell not containing $P$, we do a more elaborate modification of the arrangement described below.
We first need some terminology. Let $R$ be a digon bounded by the sweep curve $\gamma$ and another curve $\mu$ with vertices $b_{1}$ and $b_{2}$. A proper chord $\lambda$ of $R$ is an arc of another curve $\delta$ which intersects the boundary of $R$ exactly once on $\mu$ and once on $\gamma$. We say that a sequence of proper chords $\lambda_{1},\cdots,\lambda_{k}$ are parallel if for $i=1,\ldots,k-1$, the region of $R$ between consecutive chords $\lambda_{i}$ and $\lambda_{i+1}$ is a four-sided cell in the arrangement that does not contain $P$. For each $i$, let $\delta_{i}$ be the curve containing $\lambda_{i}$ and let $\Delta_{\mu}$ denote the sequence $\Delta_{\mu}=(\delta_{1},\ldots,\delta_{k})$.
Lemma 7.6.
It is possible to modify the curves in $\Delta_{\mu}=(\delta_{1},\cdots,\delta_{k})$, possibly discarding some of them, so that i) the modified arrangement is as simple, or simpler than $\Gamma$, ii) the new digon/triangle cells created in the modified arrangement in $\tilde{\gamma}$ have either $b_{1}$ or $b_{2}$ as a vertex, and iii) there are no sweepable digon/triangle cells in $\tilde{\gamma}$ on $\mu$ other than those having $b_{1}$ or $b_{2}$ as one of the corners.
Proof 7.7.
If there are no digon/triangle cells on $\beta$, then there is nothing to do. Suppose there is a digon cell on $\beta$ formed by curve $\delta_{i}$. Then, $\Gamma\setminus\{\delta_{i}\}$ is a simpler arrangement with fewer digon/triangle cells on $\beta$. Since $\delta_{i}$ forms a digon cell on $\beta$, it forms a side of a cell $C$ in $\tilde{\gamma}\setminus R$ on $\beta$. Removing $\delta_{i}$ therefore decreases the number of sides in $C$ by 1, and does not affect any other cell in $\tilde{\gamma}\setminus R$. Hence, only $C$ can become a triangle. If $\delta_{i}$ and $\delta_{i+1}$ form a triangle cell $T$ on $\beta$, let $p$ be the vertex defined by $\delta_{i}$ and $\delta_{i+1}$ that forms a vertex of $T$. Then, the triangle formed by $\delta_{i}$ and $\delta_{i+1}$ on $\gamma$ is a minimal triangle with base $\gamma$. Hence, by Corollary 5.4 we obtain a simpler 2-intersecting arrangement. Since $T$ is intersected by $\beta$, by Lemma 7, the only cells with reference points $p_{\delta_{i}}$ or $p_{\delta_{i+1}}$ can become triangle cells. Therefore, the number of triangle cells on $\beta$ increases by 2, and no new triangle cells are created. If $\delta_{i}$ is not the first or last curve in the sequence $\Delta$, then consider the two cells with reference points $p_{\delta_{i}}$ or $p_{\delta_{i+1}}$. If either cell is a cell on $\gamma$, then they have at least 4 sides, as they have one of $\delta_{i}$ or $\delta_{i+1}$, $\beta$, and one of $\delta_{i^{\prime}}$ as three sides, besides $\gamma$ as sides, where $i^{\prime}<i$ or $i^{\prime}>i+1$.
At each iteration, we obtain a simpler arrangement where either the number of curves in $\Delta$ decreases, or the number of intersections between the curves $\Delta$ decreases. Therefore, after a finite number of iterations, there are no triangle cells or digon cells on $\beta$ except possibly the ones with $b_{1}$ or $b_{2}$ as a vertex. Figures 15 show the modification of the curves in $\Delta_{\mu}$.
(a) Parallel curves on $\mu$.
(b) Triangle bypassing on $\delta_{2},\delta_{3}$, and removing $\delta_{5}$.
Figure 15: The figure shows the parallel curves on $\beta$, and two steps in the modification of the parallel curves in $\Delta$. Triangle bypassing of the triangle cell on $\mu$ formed by $\delta_{2}$ and $\delta_{3}$, and removing the digon cell formed by $\delta_{5}$. The modified curves in $\Delta$ are shown as dotted curves. Observe that $\delta_{4}$ and $\beta$ form a triangle cell on $\gamma$.
Handling triangle cells $C_{\alpha}$ and/or $C_{\beta}$: Suppose that $C_{\alpha}$ is a triangle cell $T$ that does not contain $P$, defined by the curves $\alpha,\gamma$ and $\delta_{1}$. Let $R$ be the digon defined by $\alpha$ and $\gamma$ that contains $T$. Without loss of generality let $a_{1},c_{1},d_{1}$ denote the three vertices of $T$, where $c_{1}$ and $d_{1}$ are the intersection points of $\delta_{1}$ with $\gamma$ and $\alpha$, respectively. Let $\Delta_{\alpha}=(\delta_{1},\ldots,\delta_{k})$ be a maximal sequence of parallel curves in $R$ starting with $\delta_{1}$. Since Let $d_{k}$ be the intersection point of $\delta_{k}$ and $\gamma$ in $R$. Since $\Gamma$ is a non-sweepable arrangement, either the arc of $\gamma$ between $d_{k}$ to $b_{2}$ does not bound a triangle cell, or the triangle cell contains $P$, and is therefore non-sweepable. Hence, the arc of $\gamma$ between $d_{k}$ to $b_{2}$ is intersected by a curve in $\Gamma\setminus\{\beta,\delta_{k}\}$. Similarly, since $\Gamma$ is non-sweepable, the arc of $\gamma$ between $a_{1}$ and $b_{1}$ is intersected by a curve in $\Gamma\setminus\{\beta,\gamma\}$, or is a triangle cell containing $P$. We first apply Lemma 7.6 with $\mu=\alpha$, $R$, and $\Delta_{\alpha}$. Let $\Delta_{\alpha}=(\delta^{\prime}_{1},\ldots,\delta^{\prime}_{\ell})$ denote the sequence of parallel curves after the application of Lemma 7.6. By Lemma 7.6, the possible sweepable triangle cells in the new arrangement are those in $\tilde{\gamma}$ that contain $a_{1}$ or $a_{2}$ as a vertex, i.e., the cells containing reference points $x_{1},y_{1},x_{2}$ or $y_{2}$ (See Figure 15(a)). The cell containing reference point $y_{1}$ cannot be a sweepable cell it either contains $P$, or there is at least one more curve other than $\gamma$ defining this cell that does not intersect $\alpha$ in $\tilde{\gamma}$. Similarly, the cell containing the reference point $x_{2}$ cannot be a sweepable cell, as this would imply that the cell in $\Gamma$ containing $x_{2}$ was already sweepable. Therefore, only the cells containing reference points $x_{1}$ or $y_{2}$ can be sweepable triangle cells. Observe that $\alpha,\gamma$ and $\delta^{\prime}_{\ell}$ form a minimal triangle with base $\delta^{\prime}_{\ell}$. By Corollary 5.4, bypassing this minimal triangle yields a simpler non-sweepable arrangement of 2-intersecting curves such that the two cells with reference points $x_{1}$ and $y_{1}$ merge to create a non-sweepable cell, and the cell containing reference point $y_{2}$ becomes a 4-sided cell. See Figure 15(b). Thus, we obtain a simpler non-sweepable arrangement. If $C_{\beta}$ is a triangle cell, we treat it similarly. Thus, in all cases, we obtain a simpler non-sweepable arrangement, contradicting the assumption that $\Gamma$ was a simplest non-sweepable arrangement. Hence, we can apply one of the sweep operations, and by Corollary 3, each sweep operation preserves the 2-intersecting property of the curves.
(a) $\alpha$ is a side of a triangle on bypassing $L$. The grey arcs are the parallel arcs.
(b) Bypass the minimal triangle with base $\delta^{\prime}_{\ell}$. The curves $\gamma$ and $\alpha$ intersect outside $\delta^{\prime}_{\ell}$.
Figure 16: If $\alpha$ defines a triangle on bypassing $L$, then we consider the minimal triangle defined by $\gamma,\alpha$ and $\delta$. Bypassing this minimal triangle yields a smaller non-sweepable arrangement. The figure shows that the cell with reference point $y_{2}$ could be a triangle cell on applying Lemma 7.6, but becomes 4-sided on bypassing a minimal triangle defined by $\delta^{\prime}_{\ell},\alpha$ and $\gamma$.
8 Conclusion
We gave a simpler proof of the result of Snoeyink and Hershberger [23], in the process developing conceptual tools for modifying arrangements of regions. We also showed that focusing on the more elementary condition of “non-piercing" leads to a significantly shorter and conceptually simpler proof. In Section 3, we showed that this simpler condition suffices, or yields more general results.
Snoeyink and Hershberger showed that their result does not extend to $k$-intersecting families of curves for $k\geq 3$. A natural question is whether it can be extended for $k$-intersecting non-piercing families of curves (called $k$-admissible in the literature). Unfortunately, this is not possible either - Figure 17 shows a counter-example with $5$ curves which are $4$-admissible but any process of shrinking the sweep $\gamma$ results in $\gamma$ intersecting one of the curves in $6$ points in between. Another basic result in the paper of Snoeyink and Hershberger is an extension of the Levi extension theorem. In particular, the authors show that for any set of 2-intersecting curves and 2 points not on the same curve, we can draw another curve through the two points so that the arrangement remains 2-intersecting. This theorem also extends to non-piercing regions. In particular, given a set of non-piercing curves and 3 points not all on the same curve, we can add a new curve that passes through the 3 points and the arrangement remains non-piercing. This result will appear in a companion paper.
(a) The only possible move is to pass one of the highlighted vertices
(b) Sweep curve $\gamma$ becomes $6$-admissible with the curve marked in black
Figure 17: Arrangement of $4$-intersecting non-piercing curves that can not be swept with the curve $\gamma$ maintaining $4$-admissibility.